# Defining subsystems in Hilbert spaces with non-Euclidean metric

### v1

@@ III Tomography and the no-signaling principle
A tomographically complete set of observables, given by $\{A_{i}\}\in\mathcal{A}$, satisfies the condition: $\omega_{1}(A_{i})=\omega_{2}(A_{i})\iff\omega_{1}=\omega_{2}$ for all states $\omega_{1}$ and $\omega_{2}$. Therefore, such a set completely characterizes any given state $\omega$. We will show that the corresponding density matrices, $\bar{\rho}\in\mathcal{H}_{G}$, and $\rho_{\eta}\in\mathcal{H}$ are tomographically equivalent i.e., based on the distribution of $\omega$ on a tomographically complete set of observables, the experimentalist can construct either of the states. The density matrix for the state, with respect to the representation $\pi$, is constructed through a sampling operator $W$[31] which is given as
@@
Here $\mu$ is a probability measure on the set of observables $\pi(\mathcal{A})=\mathcal{B}(\mathcal{H})$ and $\langle\mathcal{O}_{i}|\,\rho\rangle=tr(O_{i}^{\dagger}\rho)=\omega(A_{i})$ is the Hilbert-Schmidt inner product between the operators $\mathcal{O}_{i}=\pi(A_{i})$ and $\rho$. Faithfully reconstructing the state requires that the operator $W$ be full rank, which is a constraint on the measure $\mu$. It is said to be a tomographically complete measure if $W$ is full rank. Note that given a tomographically complete set of observables $\{\mathcal{O}_{i}\}$ in the Hilbert space $\mathcal{H}$, the set of observables $\{\bar{\mathcal{O}}_{i}\}$, where $\bar{\mathcal{O}_{i}}=\eta^{-1}\mathcal{O}_{i}\eta$, is a tomographically complete set in $\mathcal{H}_{G}$. Furthermore, since the norm in the two Hilbert spaces is related by $||\mathcal{O}||=||\bar{\mathcal{O}}||_{G}$, the measure in the space $\mathcal{B}(\mathcal{H}_{G})$ defined by $d\mu_{G}(\bar{\mathcal{O}})=d\mu(\mathcal{O})$ is also a probability measure. Combining these observations we can construct the sampling operator in $\mathcal{B}(\mathcal{H}_{G})$: $\bar{W}=\eta^{-1}W\eta$. Therefore, it becomes clear that if the measure $\mu$ gives a tomographically complete set of observables in $\mathcal{B}(\mathcal{H})$, then $\mu_{G}$ is tomographically complete. Furthermore, if $\mu$ is a tight frame (i.e. $W=\mathbb{1}$), then $\mu_{G}$ is also a tight frame.
@@
In an experiment, one has the data in terms of the choice of the observable to be measured (e.g. $A_{i}$) and the resulting value of the measurement $m$. The connection between this data and the equation (2) is given by Born’s rule which tells us that the inner product $\langle\mathcal{O}_{i}|\rho\rangle$ is the expectation value of the measurement results of the observable $A_{i}$: $\sum_{m}mp(m|A_{i})=\langle\mathcal{O}_{i}|\rho\rangle=\omega(A_{i})$. From our discussion in the previous sections, we know that this quantity is independent of the chosen cyclic representation of the system $(\mathcal{A},\omega)$. Indeed, in the Hilbert space with the metric $G$, this value is equal to the expectation value of the observable $\bar{\mathcal{O}}$, i.e. $\langle\mathcal{O}|\,\rho\rangle=\langle\bar{\mathcal{O}}|\,\bar{\rho}\rangle$.
@@
Take the example of the spin 1/2 system. The set $\{\mathbb{1},\vec{\sigma}\}$ forms a tight frame in the trivial metric Hilbert space $\mathcal{H}=\mathbb{C}^{2}$. It is also the minimal group that induces an irreducible representation on $\mathcal{H}$[32]. The tight frame consists of measurement of the spin in three orthogonal directions since the Pauli matrices form the orthogonal basis for the space of observables in two dimensions: $tr(\sigma_{i}\sigma_{j})=2\delta_{ij}$. A general observable is therefore given by $\vec{\sigma}.\vec{n}$ and is associated with a direction $\vec{n}$ in 3-dimensional space. Let $p(m|\hat{n}_{i})$ be the probability of getting the outcome $m$, given the spin was measured in the $\hat{n}_{i}$ direction, corresponding to the measurement of the $\sigma_{i}$ observable. To make this measurement we use a Stern-Gerlach set-up where the particles of a specific velocity pass through the magnetic field $\vec{B}=B\hat{n}_{i}^{\perp}$ (see D) and then through a magnetic field in $z$-direction which splits the beam in two directions corresponding to the values of $m\in\{\pm\dfrac{1}{2}\}$. By controlling the direction and magnitude of the magnetic field $\vec{B}$, one produces the experimental data and constructs the density matrix: $\rho=\mathbb{1}/2+\sum_{m,i}mp(m|\hat{n}_{i})\sigma_{i}$. Note that while associating the operator $\sigma_{i}$ with the direction $\hat{n}_{i}$, the experimentalist makes two choices. The first is the choice of a tight frame from the family of unitarily equivalent tight frames. Since the set of operators $\{\mathbb{1},U\sigma_{i}U^{\dagger}\}$ also forms a tight frame in $\mathcal{H}$, the density matrix thus constructed is fixed only up to a unitary equivalence. The second choice is the metric space in which the operators $\{\sigma_{i}\}$ and their unitary equivalents are observables. One can equivalently choose the set of operators $\{\bar{\sigma}_{i}=\eta^{-1}\sigma_{i}\eta\}$ which are the observables in the metric space $\mathcal{H}_{G}$ with metric $G=\eta^{2}$. They also form a tight frame with $tr(\bar{\sigma}^{*}_{i}\bar{\sigma}_{j})=2\delta_{ij}$. With this choice of the metric space, the resultant matrix $\bar{\rho}=\mathbb{1}/2+\sum_{m,i}mp(m|n_{i})\bar{\sigma}_{i}$ is a density matrix in $\mathcal{H}_{G}$.
@@
Consider the non-trivial metric representation of the system of two particles in $\mathcal{H}_{G}$ as the state $\bar{\rho}_{12}$. Alice, still in possession of the data $\{m,p(m,\hat{n}^{1}_{i})\}$, might be aware of the global metric structure $G$, but due to the non-separability of the metric, she does not have a notion of local observables. Therefore, she cannot construct a density matrix for particle 1 alone which can be seen as a reduced state of $\bar{\rho}_{12}$. From the global point of view, we have $\sum m_{1}p(m_{1}|\hat{n}^{1}_{i})=tr(\bar{\rho}_{12}\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta)$. This is because we have $tr(\bar{\sigma}_{j}^{*}\bar{\sigma}_{j})=2\delta_{ij}$, therefore, these measurements correspond to measurements in orthogonal directions. Therefore, Alice’s data also corresponds to the set of observables $\{\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta\}$, which forms a sub-algebra.
This discussion motivates us to make the following remark:
Remark: Given a density matrix $\bar{\rho}_{12}$ in $\mathcal{H}_{G}$, the density matrix $\rho_{1}=tr_{2}(\eta\bar{\rho}_{12}\eta^{-1}$) represents a well-defined sub-system of the system represented by $\bar{\rho}_{12}$.
@@
The second equality comes from the no-signaling principle in trivial metric space: $tr_{2}(\eta\pi_{G}(B)\bar{\rho}_{12}\eta^{-1})=tr_{2}(\mathbb{1}\otimes\mathcal{O}_{B}\rho_{12})=tr_{2}(\rho_{12})$. Since we have already established that Alice’s subsystem is faithfully represented by the state $\rho_{1}$, we can conclude that the no-signaling principle for a bi-partition, as defined in the definition 1, is obeyed in every metric space. In the literature [15], one typically shows that the no-signaling principle is satisfied when the metric is separable $G=G_{1}\otimes G_{2}$, which is only a specific case of the more general principle that we have exhibited here through equation (3).
@@ IV Conclusion
The axiom of the tensor product decomposition of composite systems restricts the way we understand a subsystem. In a vast family of quantum systems, such a decomposition is not possible, and consequently, a partial trace operation is not well defined. This issue arises in the non-trivial metric space representation of composite systems, which forms a natural description of the systems undergoing balanced gain and loss under the pseudo-Hermitian Hamiltonian. In this work, we have used algebraic tools to provide an operational definition of subsystems in such Hilbert spaces with non-trivial metric structures. We exhibit the fact that given an algebra and state over it, one can equivalently construct a Hermitian or a pseudo-Hermitian cyclic representation, which are unitarily equivalent if we use the term unitary in its most general sense. While given a pseudo-Hermitian Hamiltonian, one can choose from a family of metric spaces, we show that different choices of the metric operator can correspond to different choices of the system decomposition. Since choosing a metric is equivalent to choosing a reference frame for the system, it makes sense that choosing two different metric representations changes the subs-system properties. Furthermore, the choice of sub-system decomposition is equivalent to a choice of an equivalence class of metric spaces. Finally, we show that from an experimentalist’s point of view, local tomography on a sub-system undergoing pseudo-Hermitian evolution creates no paradoxes, and allows one to construct a local description of the subsystem in trivial metric space. The no-signaling principle, which so far has been restricted to tensor product decomposable Hilbert spaces, can be consistently shown to hold good for an algebraically defined sub-system, even when the Hilbert spaces do not have a genuine tensor product structure.
@@
After the mathematical description of pseudo-Hermitian quantum mechanics as a trivial extension of normal quantum mechanics was established, most of the works dealing with composite systems in this field have been wary of using the non-separable metric structure due to the ill-defined partial trace operation in such scenarios. From this work, it is clear that such metric operators are as physically meaningful as the separable ones, and there is no clear physical reasoning for choosing a separable metric operator over the non-separable operators to define the subsystems. We hope that this work will be helpful in a better understanding of composite quantum systems evolving under the pseudo-Hermitian Hamiltonian.
@@
Acknowledgments: H.B. would like to thank A P Balachandran for useful discussions and Namit Anand for his valuable comments on the early version of this draft.
Appendix A Local unitary condition
Here we show that for an irreducible set of operators $\{\mathcal{O}_{i}\}$, if the unitarily equivalent set is $\{\mathcal{O}^{\prime}_{i}\}$, s.t. $U\mathcal{O}_{i}U^{\dagger}=\mathcal{O}_{i}^{\prime}$, then this unitary is fixed up to a phase factor.
Proof.
Let $U_{1}$ and $U_{2}$ be two unitary operators s.t. $\forall i$
$U_{\alpha}\mathcal{O}_{i}U_{\alpha}^{\dagger}=\mathcal{O}_{i}^{\prime},\ \alpha\in\{1,2\}$ (4)
Then $X=U_{1}U_{2}^{\dagger}$ is an operator which commutes with all the elements of $\{\mathcal{O}_{i}\}$. If $\{\mathcal{O}_{i}\}$ is an irreducible set of operators, we must have $X=\mathbb{1}$ up to a phase. ∎
Therefore, for a set of operators $\{\mathcal{O}_{A}\otimes\mathcal{O}_{B}\}$ if we have $U\mathcal{O}_{A}\otimes\mathcal{O}_{B}U^{\dagger}=\mathcal{O}^{\prime}_{A}\otimes\mathcal{O}^{\prime}_{B}$ then $U=U_{A}\otimes U_{B}$ up to a phase.
Appendix B Lindbladian dynamics in $\mathcal{H}_{G}$
Given the Lindblad dynamics
$\dot{\rho}(t)=-i[H,\rho(t)]+\sum_{j}-\dfrac{1}{2}\{L_{j}^{\dagger}L_{j},\rho(t)\}+L_{j}\rho L_{j}^{\dagger}$ (6)
Define $H_{e}=H-i\sum_{i}L_{j}^{\dagger}L_{j}$ so that this can be re-written as
$\dot{\rho}(t)=-i(H_{e}\rho(t)-\rho(t)H_{e}^{\dagger})+\sum_{j}L_{j}\rho(t)L_{j}^{\dagger}$ (7)
@@ Appendix B Lindbladian dynamics in $\mathcal{H}_{G}$
Define $M_{j}=L_{j}\sqrt{dt}$ and $M_{0}=\mathbb{1}-iH_{e}dt$, then the following infinitesimal CP map corresponds to the above Lindblad equation
$\rho(t+dt)=M_{0}\rho(t)M_{0}^{\dagger}+\sum_{j}M_{j}\rho(t)M_{j}^{\dagger}$ (8)
@@
The probability of the jump in this infinitesimal time is given by $dp=\sum_{j}tr(L_{j}^{\dagger}L\rho(t))dt$ and the probability of no jump is $p_{0}=1-dp=tr(M_{0}^{\dagger}M_{0}\rho(t))$.
@@
Now the jump probability within the time interval $t$ to $t+dt$ is $dp=\sum_{j}tr(M_{j}^{*}M_{j}\bar{\rho})=\sum_{j}tr(L_{j}^{\dagger}\eta^{2}L_{j}\rho)$ and the probability of no jump is $p_{0}=tr(M_{0}^{*}M_{0}\bar{\rho})$. However, $M_{0}^{*}M_{0}=\mathbb{1}$, therefore $p_{0}=tr(\bar{\rho})$. One can always choose $\eta^{2}$ s.t. $tr(\bar{\rho})=1$. The pseudo-Hermiticity condition tells us $\sum_{j}tr(L_{j}^{\dagger}L_{j}\eta^{2})=0$. From here we can conclude $\sum_{j}tr(L_{j}^{\dagger}\eta^{2}L_{j}\rho)=0$
Appendix C Ambiguities in entropy of the state

### v2

@@ III Tomography and the no-signaling principle
A tomographically complete set of observables, given by $\{A_{i}\}\subset\mathcal{A}$, satisfies the condition: $\omega_{1}(A_{i})=\omega_{2}(A_{i})\iff\omega_{1}=\omega_{2}$ for all states $\omega_{1}$ and $\omega_{2}$ on $\mathcal{A}$. Therefore, such a set completely characterizes any given state $\omega$. We will show that the corresponding density matrices, $\bar{\rho}\in\mathcal{H}_{G}$, and $\rho_{\eta}\in\mathcal{H}$ are tomographically equivalent i.e., based on the distribution of $\omega$ on a tomographically complete set of observables, the experimentalist can construct either of the states. The density matrix for the state, under the representation $(\mathcal{H},\pi)$, is constructed through a sampling operator $W$[38] which is given as
@@
Here $\mu$ is a probability measure on the set of observables $\pi(\mathcal{A})=\mathcal{B}(\mathcal{H})$ and $\langle\mathcal{O}_{i}|\,\rho\rangle=tr(O_{i}^{\dagger}\rho)=\omega(A_{i})$ is the Hilbert-Schmidt inner product between the operators $\mathcal{O}_{i}=\pi(A_{i})$ and $\rho$. Faithfully reconstructing the state requires that the operator $W$ be full rank, which is a constraint on the measure $\mu$. It is said to be a tomographically complete measure if $W$ is full rank. Note that given a tomographically complete set of observables $\{\mathcal{O}_{i}\}$ in the Hilbert space $\mathcal{H}$, the set of observables $\{\bar{\mathcal{O}}_{i}\}$, where $\bar{\mathcal{O}_{i}}=\eta^{-1}\mathcal{O}_{i}\eta$, is a tomographically complete set in $\mathcal{H}_{G}$. Furthermore, since the norm in the two Hilbert spaces is related by $||\mathcal{O}||=||\bar{\mathcal{O}}||_{G}$, the measure in the space $\mathcal{B}(\mathcal{H}_{G})$, defined by $d\mu_{G}(\bar{\mathcal{O}})=d\mu(\mathcal{O})$, is also a probability measure. Combining these observations we can construct the sampling operator in $\mathcal{B}(\mathcal{H}_{G})$: $\overline{W}=\eta^{-1}W\eta$. Therefore, it becomes clear that if the measure $\mu$ gives a tomographically complete set of observables in $\mathcal{B}(\mathcal{H})$, then $\mu_{G}$ is tomographically complete. Furthermore, if $\mu$ is a tight frame (i.e. $W=\mathbb{1}$), then $\mu_{G}$ is also a tight frame.
@@
In an experiment, the knowledge of a state is in terms of the statistical data of the observable (e.g. $A_{i}$) and the probability of the value of the measurement being $m$, denoted by $\{m,p(m|A_{i})\}$. The connection between this data and the equation (3) is given by Born’s rule which tells us that the inner product $\langle\pi(A_{i})|\rho)\rangle$ is the expectation value of the measurement results of the observable $A_{i}$: $\sum_{m}mp(m|A_{i})=\langle\pi(A_{i})|\rho\rangle=\omega(A_{i})$. From our discussion in the previous sections, we know that this quantity is independent of the chosen cyclic representation of the system $(\mathcal{A},\omega)$ i.e. $\langle\mathcal{O}|\,\rho\rangle=\langle\bar{\mathcal{O}}|\,\bar{\rho}\rangle$.
@@
Take the example of the spin 1/2 system. The set $\{\mathbb{1},\vec{\sigma}\}$ forms a tight frame in the trivial metric Hilbert space $\mathcal{H}=\mathbb{C}^{2}$. This is also the minimal group that induces an irreducible representation on $\mathcal{H}$[39]. The tight frame consists of measurement of the spin in three orthogonal directions since the Pauli matrices form the orthogonal basis for the space of observables in two dimensions: $tr(\sigma_{i}\sigma_{j})=2\delta_{ij}$. A general observable is therefore given by $\vec{\sigma}.\hat{n}$ and is associated with a direction $\hat{n}$ in 3-dimensional space. Let $p(m|\hat{n}_{i})$ be the probability of getting the outcome $m$, given the spin was measured in the $\hat{n}_{i}$ direction, corresponding to the measurement of the $\sigma_{i}$ observable. To make this measurement we use a Stern-Gerlach set-up where the particles of a specific velocity pass through the magnetic field $\vec{B}=B\hat{n}_{i}^{\perp}$ (see the appendix C for details) and then through a magnetic field in $z$-direction which splits the beam in two parts corresponding to the values of $m\in\{\pm\dfrac{1}{2}\}$. By controlling the direction and magnitude of the magnetic field $\vec{B}$, one produces the experimental data and constructs the density matrix: $\rho=\mathbb{1}/2+\sum_{m,i}mp(m|\hat{n}_{i})\sigma_{i}$. Note that while associating the operator $\sigma_{i}$ with the direction $\hat{n}_{i}$, the experimentalist makes two choices. The first is the choice of a tight frame from the family of unitarily equivalent tight frames. Since the set of operators $\{\mathbb{1},U\sigma_{i}U^{\dagger}\}$ also forms a tight frame in $\mathcal{H}$, the density matrix thus constructed is fixed only up to a unitary equivalence. The second choice is the metric space in which the operators $\{\sigma_{i}\}$ and their unitary equivalents are observables. One can equivalently choose the set of operators $\{\bar{\sigma}_{i}=\eta^{-1}\sigma_{i}\eta\}$ which are the observables in the metric space $\mathcal{H}_{G}$ with metric $G=\eta^{2}$. They also form a tight frame with $tr(\bar{\sigma}^{*}_{i}\bar{\sigma}_{j})=2\delta_{ij}$. With this choice of the metric space, the resultant matrix $\bar{\rho}=\mathbb{1}/2+\sum_{m,i}mp(m|n_{i})\bar{\sigma}_{i}$ is a density matrix in $\mathcal{H}_{G}$.
@@
Consider the non-trivial metric representation of the system of two particles in $\mathcal{H}_{G}$ as the state $\bar{\rho}_{12}$. Alice, still in possession of the data $\{m,p(m,\hat{n}^{1}_{i})\}$, might be aware of the global metric structure $G$, but due to the non-separability of the metric, she does not have a notion of local observables. Therefore, she cannot construct a density matrix for particle 1 alone which can be seen as a reduced state of $\bar{\rho}_{12}$. Nevertheless, from the global point of view, we have $\sum m_{1}p(m_{1}|\hat{n}^{1}_{i})=tr(\bar{\rho}_{12}\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta)$. This is because we have $tr(\bar{\sigma}_{j}^{*}\bar{\sigma}_{j})=2\delta_{ij}$, which means that these measurements correspond to measurements in orthogonal directions. Therefore, Alice’s data also corresponds to the set of observables $\{\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta\}$, which forms a sub-algebra. With $\{\eta^{-1}(\mathbb{1}\otimes\sigma_{i})\eta\}$ forming the complementary sub-algebra together generating the total $C^{*}$ algebra $\mathcal{B}(\mathcal{H}_{G})$, Alice’s data corresponds to a well-defined subsystem of $\mathcal{H}_{G}$. But as we have established, the data $\{m,p(m|\hat{n}_{i}^{1})\}$ is faithfully represented by $\rho_{1}\in\mathcal{H}_{1}$, therefore, it captures a well-defined subsystem of $\mathcal{H}_{G}$. These arguments can be generalized for an arbitrary finite-dimensional system, and therefore, we have the following theorem:
Theorem 2.
Given a density matrix $\bar{\rho}_{12}$ in $\mathcal{H}_{G}$, the density matrix $\rho_{1}=tr_{2}(\eta\bar{\rho}_{12}\eta^{-1}$) represents a well-defined sub-system of the system represented by $\bar{\rho}_{12}$.
@@
The second equality comes from the no-signaling principle in trivial metric space: $tr_{2}(\eta\pi_{G}(B)\bar{\rho}_{12}\eta^{-1})=tr_{2}(\mathbb{1}\otimes\mathcal{O}_{B}\rho_{12})=tr_{2}(\rho_{12})$. Since we have already established that the state $\rho_{1}$ faithfully represents Alice’s subsystem, we can conclude that the no-signaling principle for a bi-partition, as given in the definition 1, is obeyed in every metric space. In the literature [15], one typically shows that the no-signaling principle is satisfied when the metric is separable $G=G_{1}\otimes G_{2}$, which is only a specific case of the more general principle that we have exhibited here through equation (4).
@@ IV Conclusion
The axiom of the tensor product decomposition of the Hilbert space of composite quantum systems restricts how we understand a subsystem. In a vast family of quantum systems, such a decomposition is not possible, and consequently, a partial trace operation is not well defined. This issue arises in the non-trivial metric space representation of composite systems, which forms a natural description of the systems undergoing balanced gain and loss under the pseudo-Hermitian Hamiltonian. In this work, we have used algebraic tools to provide an operational definition of subsystems in such Hilbert spaces with non-trivial metric structures. We show that given an algebra and a state over it, one can equivalently construct a Hermitian or a pseudo-Hermitian cyclic representation, which are unitarily equivalent if we use the term unitary in its most general sense. While given a pseudo-Hermitian Hamiltonian, one can choose from a family of Hamiltonian compatible metric spaces, we show that different choices of the metric operator can correspond to choosing different system decompositions. Since choosing a metric is equivalent to choosing a reference frame for the system, it makes sense that choosing two different metric representations changes the subs-system properties. Furthermore, the choice of sub-system decomposition is equivalent to a choice of an equivalence class of metric spaces. Finally, we show that from an experimentalist’s point of view, local tomography on a sub-system undergoing pseudo-Hermitian evolution creates no paradoxes, and allows one to construct a local description of the subsystem in trivial metric space. The no-signaling principle, which so far has been restricted to tensor product decomposable Hilbert spaces, can be consistently shown to hold good for an algebraically defined sub-system, even when the Hilbert spaces do not have a genuine tensor product structure.
@@
After the mathematical description of pseudo-Hermitian quantum mechanics as a trivial extension of normal quantum mechanics was established, most studies dealing with composite systems in this field have been wary of using the non-separable metric structure due to the ill-defined partial trace operation in such scenarios. From this work, it is clear that such metric operators are as physically meaningful as the separable ones, and there is no clear physical reasoning for choosing a separable metric operator over the non-separable operators to define the subsystems. We hope that this work will be helpful in a better understanding of composite quantum systems evolving under the $\mathcal{P}\mathcal{T}$-symmetric non-Hermitian Hamiltonians.
@@
Acknowledgments: H.B. thanks A P Balachandran for useful discussions and Namit Anand for his valuable comments on the draft.
Appendix A Lindbladian dynamics in $\mathcal{H}_{G}$
Given the Lindblad dynamics of the state $\rho$
$\dot{\rho}(t)=-i[H,\rho(t)]+\sum_{j}-\dfrac{1}{2}\{L_{j}^{\dagger}L_{j},\rho(t)\}+L_{j}\rho L_{j}^{\dagger},$ (5)
define $H_{e}=H-i\sum_{i}L_{j}^{\dagger}L_{j}$ (effective non-Hermitian Hamiltonian) so that the above equation is can be recast in the following form
$\dot{\rho}(t)=-i(H_{e}\rho(t)-\rho(t)H_{e}^{\dagger})+\sum_{j}L_{j}\rho(t)L_{j}^{\dagger}.$ (6)
@@ Appendix A Lindbladian dynamics in $\mathcal{H}_{G}$
This shows that the Lindblad equation can be written as an evolution under a non-Hermitian Hamiltonian and an extra term called a “jump term”. Define $M_{j}=L_{j}\sqrt{dt}$ and $M_{0}=\mathbb{1}-iH_{e}dt$, where $dt$ is an infinitesimal amount of time. Then the following infinitesimal CP map corresponds to the above Lindblad equation:
$\Lambda_{t,t+dt}(\rho(t))=\rho(t+dt)=M_{0}\rho(t)M_{0}^{\dagger}+\sum_{j}M_{j}\rho(t)M_{j}^{\dagger}.$ (7)
@@
The probability of the jump in this infinitesimal time interval $t$ to $t+dt$ is given by $dp=\sum_{j}tr(L_{j}^{\dagger}L\rho(t))dt$ and the probability of no jump in this interval is $p_{0}=1-dp=tr(M_{0}^{\dagger}M_{0}\rho(t))$.
@@
Now the jump probability within the time interval $t$ to $t+dt$ is $dp=\sum_{j}tr(M_{j}^{*}M_{j}\bar{\rho})=\sum_{j}tr(L_{j}^{\dagger}\eta^{2}L_{j}\rho)$ and the probability of no jump is $p_{0}=tr(M_{0}^{*}M_{0}\bar{\rho})$. However, one can check that $M_{0}^{*}M_{0}=\mathbb{1}$, therefore, $p_{0}=tr(\bar{\rho})$. Since one can always choose $\eta^{2}$ s.t. $tr(\bar{\rho})=1$, the pseudo-Hermiticity condition tells us $\sum_{j}tr(L_{j}^{\dagger}L_{j}\eta^{2})=0$. From here we can conclude $\sum_{j}tr(L_{j}^{\dagger}\eta^{2}L_{j}\rho)=0$ In other words, the jump probability in the redefined Lindbladian equation vanishes if the effective Hamiltonian is Hermitian, indicating a unitary evolution.
Appendix B Ambiguities in entropy of the state